Circular No. 8614

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 T2 (CHRISTENSEN) E. J. Christensen reports his discovery of a comet on CCD images

taken with the Mt. Lemmon Survey 1.5-m reflector, the object showing a moderately condensed 7′′ coma with a faint 10′′ tail spanning p.a. 220^{o}–240^{o} on four co-added 180-s frames (discovery observation tabulated below).

2005 UT  α_{2000}  δ_{2000}  Mag.
Oct. 7.21163  0^{h}49^{m}48.^{s}96  +18^{o}35′58′′.7  19.9

The available astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2005-T66.

T = 2005 Mar. 18.933 TT  ω = 56.^{o}235
e = 0.42490  Ω = 258.152 }2000.0
q = 2.25287 AU  i = 8.762
a = 3.91735 AU n^{o} = 0.127121 P = 7.75 years

COMET P/2005 T3 (READ) M. T. Read reports his discovery of a comet on CCD mosaic images

obtained with the Spacewatch 0.9-m f/3 reflector, the object noted as diffuse with a short tail toward the southeast (discovery observation below). Follow-up observations by J. V. Scotti with the 1.8-m f/2.7 reflector on Oct. 8.41–8.48 UT show a possible very faint tail.

2005 UT  α_{2000}  δ_{2000}  Mag.
Oct. 7.33306  2^{h}00^{m}18.^{s}71  +12^{o}27′31′′.9  20.6

The available astrometry, the following very preliminary orbital elements, and an ephemeris appear on MPEC 2005-T67.

T = 2004 Nov. 20.230 TT  ω = 337.^{o}154
e = 0.26759  Ω = 27.805 }2000.0
q = 5.92218 AU  i = 5.804
a = 8.08584 AU n^{o} = 0.042866 P = 22.99 years

(809) LUNDIA A. Kryszczynska, Poznan Astronomical Observatory, and colleagues

have reported on CBET 239 that their observations indicate a binary nature for (809) and call for additional photometry and direct imaging.

2005 October 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green